OpenStudy (anonymous):

H'wat? Solve for a. 4*xy/15=88+ab-10y

OpenStudy (igreen):

Is this it? \(\dfrac{4xy}{15} = 88 + ab - 10y\)

OpenStudy (anonymous):

yeah

OpenStudy (igreen):

Okay, first we can subtract 88 to both sides: \(\dfrac{4xy}{15} - 88 = ab - 10y\) Now can you add 10y to both sides?

OpenStudy (anonymous):

I was just asking questions to give you guys something to do since no one is asking anything XD

OpenStudy (igreen):

....?

OpenStudy (igreen):

So this really isn't your problem?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

I finished all my math courses last year.

OpenStudy (igreen):

Okay, well we can add 10y to both sides: \(\dfrac{4xy}{15} - 88 + 10y = ab\) Divide 'b' to both sides: \(a =\dfrac{\dfrac{4xy}{15} - 88 + 10y}{b}\)

OpenStudy (igreen):

Okay, well thanks for the problem and medal :P

OpenStudy (igreen):

737 more medals to go..

OpenStudy (anonymous):

XD